OpenStudy (anonymous):

Can anyone walk me through the steps for this quadratic formula including what the discriminant is step by step? -x^2 - 3 +5 = 0

OpenStudy (anonymous):

i assume your equation is \[-x^2 - 3x + 5 = 0\] yes?

OpenStudy (radar):

-11 is not a perfect square, the coefficients are integers, so the solutions will be two nonreal complex values.

OpenStudy (radar):

Use the quadratic formula to solve.

OpenStudy (radar):

\[-(-3)\pm \sqrt{-3^{2}-4(-1)(5)} \over 2(-1)\]\[3\pm \sqrt{9+20}\over -2\] please note the error I made in the solution of the discriminate which should be 29 rather than -11. Thus since it is positive there will be two irrational solutions. \[3\pm \sqrt{29}\over -2\]

OpenStudy (radar):

*Correction to the earlier offending post: a= -1 b= -3 c= 5 discriminant = b^2 - 4ac that would be (-3)^2 - 4(-1)(5)= 9+20= +29 or it is positive. What does that tell you about the zeroes or solutions?

OpenStudy (radar):

That post has now been deleted. Sorry for the confusion it may have created.